OpenStudy (arihangdu):

(sqrt 7x) times ((sq rt x)-7sq rt.7) is the answer x times sq rt 7 -49 times sq rt. x?

OpenStudy (anonymous):

looks like \[\sqrt{7x}\times \sqrt{x}-7\sqrt{7}\]

OpenStudy (arihangdu):

Sqrt7x (sort X -7sqrt7)

OpenStudy (arihangdu):

Yaah

OpenStudy (anonymous):

\[\sqrt{7x}\times (\sqrt{x}-7\sqrt{7})\]?

OpenStudy (arihangdu):

Yes please

OpenStudy (arihangdu):

Please @ stile help me

OpenStudy (arihangdu):

Sort(.7x.x ) - 7sqrt(7x.x) Sort.7x^2 - 7sqrt.7^2x Xsqrt7 - 49sqrtx

OpenStudy (arihangdu):

Is this correct ?

OpenStudy (dayakar):

\[\sqrt{7x}*\sqrt{x}-\sqrt{7x}*7\sqrt{7}\]

OpenStudy (dayakar):

using distributive propertity

OpenStudy (dayakar):

\[\sqrt{7}x - 49\sqrt{x}\]

OpenStudy (arihangdu):

Thanks